A man throws a ball to maximum horizontal distance of 80 m. Calculate the maximum height reached.

Grade:

3 Answers

rahul
9 years ago
20metres
Shivam Chopra
9 years ago
We know that R = u2 * Sin 2θ / g.
As, the range is maximum, so, θ = 450
So, 80 = u2 /g.
Now, H = height reached by ball =  u2 Sin2θ / 2g.
Putting value of θ = 450 , u2 / g = 80, we get
H = 80/4  = 20m
Please approve it if you like
 
Hassnain Ali
5 years ago
A we know that Range is n times the Height of Projectile.
R=4H
As the maximum distance is 80m which is no doubt equal to 80m.So Putting values
80m=4H
80/4=H
20=H
So maximum height is 20m